Ineedhelponmymath:

In a class of 147 students, 95 are taking math (M), 73 are taking science (S), and 52 are taking both math and science. What is the probability of randomly choosing a student who is taking neither math nor science? Round your answer to the nearest tenth.

mhanifa:

Step 1. Students taking either M or S: 95 + 73 - 52 = x Step 2. Students taking neither M nor S: 147 - x = y Step 3. Required probability is: P = (y/147)*100% Find x, then y, then P, then round it to the nearest tenth